OpenStudy (anonymous):

an archer fires an arrow at 40 degrees above the horizontal. At the highest point of its trajectory, the arrow has a speed of 200 m/s. What is the arrow's initial speed?

OpenStudy (anonymous):

neglecting air drag, \[uCos \theta = 200 m/s\]\[\theta=40^{0}, \]so initial speed is \[u = 200/\cos(40^{0}) = 261.081458 m/s\]